# Efficient AI in Practice: Training and Deployment of Efficient LLMs for Industry Applications

### v1

@@ 4.3. Reasoning Task
4.3.1. Distillation Results
We investigate several distillation methods to produce a 1.5B model initialized from Qwen-2.5-1.5B-Instruct (Yang et al., 2024). The teacher models are obtained by training various sizes of Qwen-2.5 Instruct models using SFT with identical hyperparameters. Performance is primarily measured by validation loss, with lower values indicating higher accuracy (see Table 5).
@@ 4.3.1. Distillation Results
Table 5. Validation losses for various training methods and ablations. Unless explicitly specified (e.g., FKL (3B) indicates distillation using the 3B-SFT model as the teacher), the default teacher is the 7B-SFT model.
5. Deployment
5.1. RecSys use case
@@
SLM variants of the RecSys use case have been deployed to an A/B test.
@@
Serving Infrastructure For all use cases, we benchmark and serve traffic using nodes with 256 CPU cores, 2TB of host memory and 8 NVIDIA H100 GPUs. As mentioned earlier, we deploy SGLang (version 0.4.1) as the serving engine. We use tensor parallelism to concurrently use more than 1 GPU for inference. To maximize performance, we employ FP8 quantization for both weights and activations and use FlashInfer (Ye et al., 2025) as the primary attention backend. Moreover, SGLang incorporates RadixAttention, which enables prefix caching for prompts sharing common prefixes.
@@
Workloads For the RecSys workload, we follow the prompt structure outlined in Section 4.2 and utilize context lengths of 16k and 32k. Given the predictive nature of the task, the output length (i.e., the number of generated tokens) is set to 1, rendering the workload heavily reliant on the prefill phase. Consequently, optimizing the prefill stage is crucial for performance. For instance, prefix caching substantially improves prefill (and decode) times when a prompt’s key (K) and value (V) tensors for its shared prefix have already been processed. In cases where there are $k$ candidate items to be ranked for a member $m$, $k$ prompts are served. These prompts share a long prefix containing the user information and historical item interactions. Once one prompt is served, its KV tensors are cached, allowing subsequent prompts for the same member to reuse the cached data—a process we refer to as a hot prefill.
@@
Metrics We use two key metrics - time to first token (TTFT) and time per output token (TPOT). For prediction tasks, which are prefill-intensive, TTFT is the primary metric as it reflects the duration of the prefill phase. For generative tasks, both TTFT and TPOT are important. Additionally, we report the total serving throughput for various context lengths.
@@
Results In terms of quality, the models under consideration performn similarly since their AUCs are similar. For performance, we report TTFT and throughput measurements for both 16k and 32k context lengths. P99 TTFT results for a workload with 1 QPS (i.e., $m=1$) and 1 or more prompt per member (i.e., $k=1$ or more) can be found in Figure 4 (detailed p50, p99 and throughput numbers can be found in Tables 9,  10 and  11 in Appendix A.2). From the figure and the tables, we can conclude that latency drops drastically as model size becomes smaller. Serving traffic with 32k context is significantly slower than serving traffic with 16k context. Setting $k$ more than 1 doesn’t hurt latency much, because of KV caching.
@@
8B6.4B3B2.4B2.1B200200200200300300300300400400400400500500500500600600600600700700700700LLMsP99 TTFT (ms)m=1, k=1, 16km=1, k=1, 32km=1, k=4, 32k
Figure 4. P99 TTFT (ms) for various LLMs
5.2. Generative use case
Model IQM (%)
KD Model +20.29%
Table 6. Online A/B test results for reasoning task.
The reasoning task outlined in Section 4.3 was launched online for a 1% A/B test. Results are outlined in Table 6. KD, along with data changes, helped the model improve by 20.29% on an internal quality metric (IQM). We also discuss deployment lessons from a generative use case to study the effect of different quantization schemes on model inference speed and accuracy.
Serving Infrastructure Our setup is mostly similar to Section 5.1. However, in addition to using NVIDIA H100 GPUs, we also study the effect of using older NVIDIA A100 GPUs. To this end, we use the vLLM backend (version 0.6.1) for serving. We use 1 GPU for serving (TP=1).
@@ 5.2. Generative use case
Performance Results The inference speed results are reported in Table 7. Using the state-of-the-art H100 GPUs results in faster inference (both in terms of TTFT and TPOT) compared to A100 GPUs. In particular, we observe that FP8 serving with H100s leads to the smallest TTFT and TPOT. However, limiting our comparisons to A100 GPUs, we see that using INT8 (W8A8) quantization results in significant reduction of both TTFT and TPOT. On the other hand, INT4 (w4A16) quantization actually increases TTFT, while decreasing TPOT. For generative task with long output sequences, we observe that W4A16 has smaller latency. In summary, we have observed that using higher-end hardware such as H100 leads to the most speed-ups. When such hardware is not available, for generative tasks with long output sequences, W4A16 has the most speed-up, while W8A8 can be more appropriate for prefill-heavy tasks. For the sake of completeness, we present a brief comparison of quantization methods in terms of accuracy in Appendix A.1.
@@
Table 7. Comparison of different quantization methods for the Llama-3 8B model.
3B (16k)2.1B (16k)3B (32k)2.1B (32k)00555510101010Latency (ms)AttentionAllReduceMLP
@@
Figure 5. Latency breakdown of a single Transformer block for pruned and unpruned models. As context size increases, attention becomes a major bottleneck.
6. Conclusion
In this paper, we highlight the importance of training and deploying small language models (SLMs) for real-world industry use cases. We specifically drill into model distillation and compression techniques, carefully dissecting the value added by each technique for model generalization. After creating high-quality SLMs, we also detail deployment lessons learned in terms of latency and hardware efficiency. Future directions could include on-policy data-based training for more robust distillation and extending pruning to the attention parts of the transformer blocks.
Appendix A Appendix
A.1. Comparison of Quantization Methods

### v2

@@ 4 Experimental Results for Reasoning Tasks
We investigate several distillation methods to produce a 1.5B model initialized from Qwen-2.5-1.5B-Instruct Yang et al. (2024). The teacher models are obtained by training various sizes of Qwen-2.5 Instruct models using SFT with identical hyperparameters. Performance is primarily measured by validation loss, with lower values indicating higher accuracy (see Table 1).
@@
Table 1: Validation losses for various training methods and ablations. Unless explicitly specified (e.g., FKL (3B) indicates distillation using the 3B-SFT model as the teacher), the default teacher is the 7B-SFT model.
To further study the effectiveness of our distillation recipes, we apply them to other open source models. In particular, we use the Qwen3 (Yang et al., 2025) model family. On an internal reasoning task, using the same recipes discussed above, we distill a Qwen3 14B model into a Qwen3 4B student model that matches the same accuracy without any performance drops. In addition, using the OpenThoughts reasoning dataset Guha et al. (2025), we distill the Qwen3 32B reasoning model into Qwen3 8B and 4B student models. We evaluate the student models on AIME 2024 (and 2025) Habib et al. (2023) benchmarks. Using our recopies, we improve the student model’s performance by more than $20\%$ for the 8B student model and $15\%$ for the 4B student model compared to their initial base model.
Model AUC Delta (%)
8B Distilled Model -
6.4B Pruned Model (20%) + SFT -0.47%
6.4B Pruned Model (20%) + Distillation -0.06%
Table 2: Distillation vs. SFT for post-pruning retraining.
Model AUC Delta (%)
8B Model -
6.8B Pruned Model 0.0%
6.4B Pruned Model -1.33%
6.0B Pruned Model -1.72%
Table 3: Evaluation of the 8B-parameter model post-SFT and its pruned variants, focusing on MLP pruning performed in a one-shot manner (i.e., no retraining after pruning).
Model #Params AUC Delta (%)
3B (Distilled from FM) 3B -
MLP Prune + Distill 2.4B -0.12%
MLP Prune + Distill (Gradual) 2.4B 0.03%
Table 4: Comparison of one-step vs gradual pruning for a 3B model distilled from the FM.
Model #Params AUC Delta (%)
MLP Pruning 2.4B -
$\wedge$ + Attention Pruning 2.1B -1.07%
$\wedge$ + Distillation 2.1B 0.02%
Table 5: Results for attention pruning. We consider the 2.4B gradual pruning model from Table 4 as the base. The second row shows the result for one-shot attention pruning, while the last row shows the results after performing distillation.
5 Experimental Results for Predictive Tasks
5.1 Knowledge Distillation Findings
To evaluate how knowledge distillation (KD) affects model generalization, we compare it against standard supervised fine-tuning (SFT). We focus on the task performance retention of the distilled (or fine-tuned) student models relative to the original foundation model (FM). Specifically, we use Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct Dubey et al. (2024) as our student models. These models offer strong performance while remaining sufficiently compact for throughput- and latency-sensitive environments. In both KD and SFT, responses are generated from the ground-truth action-prediction labels. For KD, the per-token loss is a weighted combination of: 90% forward KL divergence between teacher and student logits, and 10% cross-entropy loss with the ground-truth labels.
Additionally, we include an extra 5% loss contribution computed over the entire sequence (including the prompt), normalized by its token count. Thus, 95% of the loss is computed solely on action prediction tokens (i.e., yes/no token), while the remaining 5% is computed over the prompt tokens. This allows the model to maintain its foundation model knowledge, helping generalization to unseen tasks.
Each KD and SFT configuration undergoes hyperparameter tuning (e.g., peak learning rate, warmup schedule, decay schedule, and weight decay). Figure 2 summarizes the results by reporting the AUC delta relative to the original FM (which achieves the best performance). The main observations are:
  • SFT: The 3B and 8B student models fine-tuned only with SFT underperform compared to the FM, which is expected given their smaller size and post-training. The performance drop for the 3B model ($1.21\%$) is larger than that of the 8B model ($0.62\%$).
  • KD: Using logit supervision from the FM consistently preserves task performance better than SFT. The 8B-KD model shows a minor AUC drop of $0.06\%$ (compared to $0.62\%$ for 8B-SFT), while the 3B-KD model (-0.15%) substantially mitigates the loss relative to 3B-SFT ($1.21\%$). These results demonstrate the effectiveness of KD for transferring knowledge.
8B-KD8B-SFT3B-KD3B-SFT1-10.5-0.500.06-0.060.62-0.620.15-0.151.21-1.21AUC Delta (%)
Figure 2: Comparison of Distillation and SFT on the Foundation Model. Knowledge distillation consistently outperforms SFT by effectively leveraging teacher supervision to preserve and enhance performance.
5.2 Post-training Compression Findings
After obtaining the distilled student models, we apply structured pruning and on-the-fly FP8 quantization to further compress the model to meet our serving latency requirements (steps outlined in Figure 1). Since the distilled models (DMs) are decoder-only transformers, our approach focuses on applying structured pruning to remove redundant MLP up/down projection neurons, as well as attention heads in each transformer layer while preserving its capabilities on in-domain ranking tasks. Since an off-the-shelf application of one-shot pruning can result in a loss in model quality (utility), we apply targeted fine-tuning after each pruning step to ensure that the model adapts to its reduced size without significant performance loss. To this end, we again leverage knowledge distillation to bridge the gap between the original and pruned models, transferring key insights and ensuring the pruned version closely aligns with the outputs of the original foundational model.
For structured pruning, we leverage the OSSCAR algorithm Meng et al. (see Section A.2 for details on methodology). We now discuss the details of various ablations that we conducted towards structured pruning.
We also explore gradual pruning Benbaki et al. (2023); Meng et al. (2024b) in which the model is pruned in multiple steps, with knowledge distillation after each pruning. Table 4 summarizes results for pruning a distilled 3B model to 2.4B (a 37.5% MLP sparsity) either in a single step or in two smaller steps (removing 1536 hidden neurons in each step, for a total of 3072). The gradual approach recovers model AUC better than the single-step approach, achieving near-lossless compression from 3B to 2.4B.
We also analyze the effect of calibration data for pruning and results are summarized in Appendix B.2
6 Deployment
6.1 Predictive use case in RecSys
@@
SLM variants of the RecSys use case have been deployed to a large-scale A/B test.
@@
Serving Infrastructure For all use cases, we benchmark and serve traffics using nodes with 256 CPU cores, 2TB of host memory and 8 NVIDIA H100 GPUs. As discussed in Appendix A.3, we deploy SGLang (version 0.4.1) as the serving engine. We use tensor parallelism to concurrently use more than 1 GPU for inference. To maximize performance, we employ FP8 quantization for both weights and activations and use FlashInfer Ye et al. (2025) as the primary attention backend. Moreover, SGLang incorporates RadixAttention, which enables prefix caching for prompts sharing common prefixes.
@@
Workloads For the RecSys workload, we follow the prompt structure outlined in Section 2 and utilize context lengths of 16k and 32k. Given the predictive nature of the task, the output length (i.e., the number of generated tokens) is set to 1, rendering the workload heavily dependent on the prefill phase. Consequently, optimizing the prefill stage is crucial for performance. For instance, prefix caching substantially improves prefill (and decode) times when a prompt’s key (K) and value (V) tensors for its shared prefix have already been processed. In cases where there are $k$ candidate items to be ranked for a member $m$, $k$ prompts are served. These prompts share a long prefix containing the user information and historical item interactions. Once one prompt is served, its KV tensors are cached, allowing subsequent prompts for the same member to reuse the cached data - the process we refer to as a hot prefill.
@@
Metrics We use two key metrics - time to first token (TTFT) and time per output token (TPOT). For prediction tasks that are prefill-intensive, TTFT is the primary metric as it reflects the duration of the prefill phase. For generative tasks, both TTFT and TPOT are important. We report the total serving throughput for various context lengths.
@@
Results In terms of quality, the models under consideration perform similarly with comparable AUCs. For performance, we report TTFT and throughput measurements for both 16k and 32k context lengths. P99 TTFT results for a workload with 1 QPS (i.e., $m=1$) and 1 or more prompt per member (i.e., $k=1$ or more) can be found in Figure 3 (detailed p50, p99 and throughput numbers can be found in Tables 8,  9 and  10 in Appendix B.3). From the results, we can conclude that latency drops drastically as model size becomes smaller. Serving traffic with 32k context is significantly slower than that with 16k context. Setting $k$ more than 1 doesn’t hurt latency much, because of KV caching.
@@
8B6.4B3B2.4B2.1B200200300300400400500500600600700700LLMsP99 TTFT (ms)m=1, k=1, 16km=1, k=1, 32km=1, k=4, 32k
Figure 3: P99 TTFT (ms) for various LLMs
6.2 Generative use case in RecSys
The reasoning task in RecSys was launched online for an 1% A/B test. Along with data changes, KD helped the model improve by 20.29% on an internal quality metric (IQM). We also discuss deployment lessons from a generative use case to study the effect of different quantization schemes on model inference speed and accuracy.
Serving Infrastructure Our setup is mostly similar to Section 6.1. However, in addition to using NVIDIA H100 GPUs, we also study the effect of using older NVIDIA A100 GPUs. To this end, we use the vLLM backend (version 0.6.1) for serving. We use 1 GPU for serving.
@@ 6.2 Generative use case in RecSys
Performance Results The inference speed results are reported in Table 6. Using the state-of-the-art H100 GPUs results in faster inference (both in terms of TTFT and TPOT) compared to A100 GPUs. In particular, we observe that FP8 serving with H100s leads to the smallest TTFT and TPOT. However, for A100 GPUs, INT4(W4A16) quantization yields the most speed-up, while INT8 (W8A8) is more appropriate for prefill-heavy tasks. For the sake of completeness, we present a brief comparison of quantization methods in terms of accuracy in Appendix B.1.
@@
Table 6: Comparison of different quantization methods for the Llama-3 8B model.
3B (16k)2.1B (16k)3B (32k)2.1B (32k)0551010Latency (ms)AttentionAllReduceMLP
@@
Figure 4: Latency breakdown of a single Transformer block for pruned and unpruned models. At longer context sizes, attention is a bottleneck.
7 Limitations
In this paper, we have studied extensively on the training and deployment of efficient SLMs for industry use cases in Recommendation Systems. We do not study any use cases beyond that; thus, our work does not cover all the techniques related to efficient LLM deployment and training. There are several limitations that we want to briefly mention here:
1. We did not include the state-of-the-art (SoTA) sparse attention techniques due to the Inference Engine compatibility issues. Recent work such as Star Attention Acharya et al. (2025) has shown promising results in offline inference speedup. We have explored SoTA sparse attention techniques and achieved >>3x of offline inference speedups. However, the SGLang inference engine currently does not support those techniques. We are working closely with the SGLang team to align with their roadmap, and hopefully to enable those features in the near future.
2. For the LLM model pruning techniques discussed in this paper, we focus only on structured pruning due to its elegant theoretical guarantees and hardware efficiency. However, recent work in unstructured pruning Jeong et al. (2025); Muñoz et al. (2024) has shown promising performance in inference speed-ups and model quality preservation. A future direction could be to combine unstructured sparsity with structured pruning to further enhance LLMs inference efficiency. To this end, one can draw inspiration from recent work on algorithmic approaches for unstructured sparsity Meng et al. (2024a); Sun et al. (2023). Other directions include exploring compression strategies adaptive to downstream parameter efficient fine-tuning Makni et al. (2025).
8 Acknowledgement
Rahul Mazumder contributed to this work while he was a consultant for LinkedIn as an Academic Scholar (in compliance with Massachusetts Institute of Technology’s outside professional activities policies).
Appendix A Methods
In this section, we detail various techniques that allow SLMs to retain strong generalization or task-specific performance, while allowing efficient serving from a latency or throughput standpoint. Specifically, we discuss training via knowledge distillation and post-training model compression. We also intersperse serving and training efficiency concerns across the entire section.
A.1 Knowledge Distillation
Knowledge distillation (KD) Hinton (2015) transfers knowledge from a larger and expressive “teacher” model to a smaller “student” model, allowing the latter to approximate teacher performance with reduced computational resources. KD can be broadly performed in two different ways (1) by leveraging the output of a teacher model to train the student Tunstall et al. (2023); Guo et al. (2025)(also known as black-box distillation) or (2) by leveraging intermediate outputs Muralidharan et al. (2024); Hinton (2015) (also known as white-box distillation). White-box distillation using the soft probabilistic outputs of the teacher is a powerful technique and helps provide richer information than hard labels used in supervised fine-tuning (SFT), helping the student generalize better, especially in tasks where smaller models struggle to discover patterns in noisy data.
$\mathcal{L}[p_{\mathbf{y}},\mathcal{D}(p\|q_{\theta})]$ $=$ (3)
  1. 1.
    Responses are drawn from $p_{\mathbf{y}}$, which may correspond to ground truth data, the teacher model (sequence-level KD)  Kim and Rush (2016), or the student model itself (on-policy KD)  Agarwal et al. (2024); Gu et al. (2024); Zhou et al. (2023). Recent advancements Xu et al. (2024); Ko et al. (2024) explore a balance between on-policy and off-policy sampling to mitigate the mismatch between student-generated responses and the teacher’s distribution while addressing inefficiencies in online student autoregressive training.
  2. 2.
  • Forward Kullback-Leibler (KL) Divergence (FKL):
    $\mathcal{D}_{\text{FKL}}\left[p(y_{t}|\mathbf{y}_{<t},\mathbf{x})\|q_{\theta}(y_{t}|\mathbf{y}_{<t},\mathbf{x})\right]$
    $=\sum_{i\in\mathcal{V}}p(i|\cdot)\log\left(\frac{p(i|\cdot)}{q_{\theta}(i|\cdot)}\right),$
  • Reverse Kullback-Leibler Divergence (RKL):
    $\mathcal{D}_{\text{RKL}}\left[p(y_{t}|\mathbf{y}_{<t},\mathbf{x})\|q_{\theta}(y_{t}|\mathbf{y}_{<t},\mathbf{x})\right]$
    $=\sum_{i\in\mathcal{V}}q_{\theta}(i|\cdot)\log\left(\frac{q_{\theta}(i|\cdot)}{p(i|\cdot)}\right),$
  • Jensen-Shannon Divergence (JSD):
    $\mathcal{D}_{JS(\beta)}\left[p(y_{t}|\mathbf{y}_{<t},\mathbf{x})\|q_{\theta}(y_{t}|\mathbf{y}_{<t},\mathbf{x})\right]$
    $=\;\beta\mathcal{D}_{\text{FKL}}\left[p\|m\right]+(1-\beta)\mathcal{D}_{\text{FKL}}\left[q_{\theta}\|m\right],$
    where $m=\beta p+(1-\beta)q_{\theta}$ is the mixture distribution.
A.2 Post-training model compression
Model compression is a widely studied area of machine learning. We specifically focus on post-training compression (PTC) techniques for improving the inference efficiency of LLMs.
In post-training compression, we apply compression to the model after training. A common compression procedure is based on pruning or quantizing the model weights of a pre-trained model but it can result in large loss in model utility due to which alternative approaches are preferred. Specifically, recent compression procedures employ a layerwise approach (Equation 4) where the utility for every layer in a model is retained (to the extent possible) by minimizing a suitable layerwise objective function based on a calibration dataset — this approach can be scaled to large models while retaining model utility, and we use this method in our work.
We describe a mathematical framework for layerwise PTC using calibration data. Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ denote the calibration data that serve as inputs to a linear layer of the model (e.g., an MLP or an attention projection). Here, $n$ is the number of tokens in the calibration dataset, and $d$ is the input dimension of the layer. For instance, in the case of an MLP down projection layer of a Transformer block, $d$ corresponds to the intermediate size of the model.
$\min_{\hat{\mathbf{W}}}\|\mathbf{X}\mathbf{W}-\mathbf{X}\hat{\mathbf{W}}\|_{F}^{2}\quad\text{subject to}\quad\hat{\mathbf{W}}\in\mathcal{Q},$ (4)
where $\mathcal{Q}\subseteq\mathbb{R}^{d\times p}$ denotes the set of feasible solutions that conform to a particular compression scheme (eg, unstructured or structured pruning, quantization, etc), In practice, the set $\mathcal{Q}$ often exhibits a discrete structure, which renders the optimization problem in (4) challenging to solve. Past work has shown that a better optimization procedure for Problem (4) generally results in better utility-compression tradeoffs Meng et al. ; Meng et al. (2024a, b); Behdin et al. (2023), which motivates the approaches we used in our work.
Here, we consider two compression techniques:
Quantization In quantization, the model weights are represented in lower precision, using a fewer number of bits. Quantization has proved to be successful in the LLM domain, to obtain compressed models with small accuracy loss Frantar et al. (2022); Xiao et al. (2023). In this work, we consider weight-only quantization, where only model weights are quantized, as well as weight and activation quantization. We study methods such as GPTQ Frantar et al. (2022) and QuantEase Behdin et al. (2023) for 4-bit weight-only quantization (aka W4A16), SmoothQuant Xiao et al. (2023) for 8-bit weight-and-activation quantization (aka W8A8), and 8-bit floating-point (FP8) quantization. Since quantization is dependent on hardware, we discuss the details of quantization-related experiments in Section 6 (deployment).
Structured Pruning Post-training neural network pruning has old roots LeCun et al. (1989); Hassibi et al. (1993) – the basic idea is to identify “redundant” model weights and set them to zero to reduce model footprint. Recently, due to increasing challenges associated with large models, advanced algorithms have been explored for pruning neural networks at the post-training stage. Some unstructured pruning methods for LLMs using layerwise reconstruction error include Meng et al. (2024a); Frantar and Alistarh (2023); Sun et al. (2023). Other pruning approaches using loss functions different from layerwise reconstruction error include approaches based on global Fisher loss (Benbaki et al., 2023; Meng et al., 2024b, and references therein), layerwise loss functions that depend upon the future layers Lucas and Mazumder (2024), and low-rank fine-tuning adapted loss functions Makni et al. (2025)[see references therein] — we leave the exploration of these approaches as interesting directions for future research.
Without any structure on the model sparsity, however, it can be difficult to realize any inference acceleration from pruning [^2] Some of the approaches discussed above apply Meng et al. (2024a); Sun et al. (2023) to semi-structured pruning such as 2:4 sparsity which is different from structured pruning.. Therefore, in this work, we pursue a structured pruning approach. In structured pruning, the goal is to obtain smaller models via removing some neurons from the model weights. In particular, we study MLP pruning, where the goal is to reduce the intermediate size of the model via removing some hidden neurons in feed-forward layers. We also study attention pruning, where we remove a certain number of attention heads from the model Meng et al. ; Kwon et al. (2022); Kurtić et al. (2024). In this paper, we use OSSCAR Meng et al. which uses a discrete optimization approach for structured pruning. OSSCAR can be scaled to the large-scale problems we consider here. We use OSSCAR as it can result in state-of-the-art performance when it comes to post-training structured pruning of LLMs and leads to the least drop in accuracy when compared to other methods.
A.3 Training and serving efficiency
Training Efficiency LLM training presents a formidable challenge due to the sheer scale of these models and the quadratic complexity of transformer architectures. Model FLOPs utilization (MFU) Chowdhery et al. (2023) is commonly used to measure GPU efficiency, making it necessary to optimize kernel operations and distributed training strategies. We have implemented Liger Kernel Hsu et al. (2025) in Triton Tillet et al. (2019), incorporating several key optimizations. First, we employ kernel fusion to reduce repetitive memory transfers between SRAM and DRAM. Next, we adopt in-place tensor modifications to avoid creating additional tensors whenever possible, thus lowering the memory footprint. We also apply chunking, which prevents the full materialization of large tensors and provides tuning flexibility while maintaining comparable performance. In combination, these optimizations reduce training time by 20% and memory usage by 60%. Additional performance-memory tradeoffs such as gradient checkpointing and CPU offloading can lead to as much as a threefold speedup.
[…]
Serving Efficiency Serving LLMs efficiently poses unique challenges due to both high computational demands and strict latency requirements. In production environments, the choice of serving frameworks is pivotal for maximizing throughput and minimizing response times. Several solutions, including vLLM, SGLag, TRT-LLM, and MLC-LLM Kwon et al. (2023); Zheng et al. (2024); NVIDIA (2024); team (2023-2025), have been proposed to address these needs. In our use cases, we extensively evaluated vLLM and SGLang. Our benchmarks revealed that SGLang is better suited to our workloads because its radix tree-based caching mechanism aligns well with our usage patterns and it integrates tightly with FlashInfer Ye et al. (2025), whose efficient attention kernels accelerate the sequence lengths and batch sizes we typically handle.
To further improve serving performance, we deploy our models on NVIDIA H100 GPUs at FP8 precision, striking a practical balance between computational efficiency and model quality. Additional details regarding our LLM serving engine configurations can be found in Section 6.
Appendix B Additional Numerical Experiments
B.1 Comparison of Quantization Methods